OpenStudy (anonymous):

Is lR^* cyclic? Try to prove your answer. Hint: Suppose k is a generator of lR^*

OpenStudy (zzr0ck3r):

wth is IR?

OpenStudy (zzr0ck3r):

do you mean \(\mathbb{R}\)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But I'm not sure what the * means.

OpenStudy (anonymous):

It's A7 chapter 11

OpenStudy (anonymous):

They have more information and a small picture.

OpenStudy (anonymous):

If k<1, then k>k^2>k^3>...

OpenStudy (anonymous):

If k>1, then k<k^2<k^3<...

OpenStudy (zzr0ck3r):

suppose to the contrary \(\mathbb{R}\) is cyclic with generator \(k\), then \(\mathbb{R}=\{0,k,k^2,k^3,....\}\) but by density for any \(k^i\ne k^j\) there exists \(r\in \mathbb{R}\) s.t. \(k^i<r<k^j\). A contradiction.

OpenStudy (zzr0ck3r):

hmm the fact that they are giving other stuff makes me think they dont want you to use density.

OpenStudy (zzr0ck3r):

Let me look...

OpenStudy (zzr0ck3r):

My wife needs me to fix something. Are you coming out tomorrow?

OpenStudy (anonymous):

Is there another way to prove this without using density?

OpenStudy (zzr0ck3r):

im sure

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

But what is r star?

OpenStudy (zzr0ck3r):

you can tell me the other ones in there, and ill know what we are doing when you get there.

OpenStudy (zzr0ck3r):

just postive R

OpenStudy (anonymous):

with what operation?

OpenStudy (zzr0ck3r):

addition I think. Look in the book in the section where it defines groups. It tells you some basic notation they use.

OpenStudy (anonymous):

My problems are A7, B4, C5, E4, and F4 Chapter 11. and on cosets.

OpenStudy (zzr0ck3r):

sweet. ok wife yelling at me. sorry. ill be there at 9

OpenStudy (anonymous):

Okay thanks Zach.

OpenStudy (zzr0ck3r):

for sure